25,54,64,93,72,87,61,39,81,47,65,53 find the means of the sample data
Answer by jim_thompson5910(35256)

To find the mean, add up the values and divide by 12. You divide by 12 since there are 12 numbers.

(25+54+64+93+72+87+61+39+81+47+65+53)/12 = 741/12 = 61.75

The mean is 61.75